# Tuning Algorithmic and Architectural Hyperparameters in Graph-Based Semi-Supervised Learning with Provable Guarantees

### v1

@@ 3 Label Propagation-based Families and Generalization Guarantees
In this section, we consider three parametric families of label propagation-based algorithms, the classical type of algorithms for semi-supervised learning. Label propagation algorithms output a soft-label score $F^{*}\in\mathbb{R}^{n\times c}$, where the $(i,j)$-th entry of $F^{*}$ represents the score of class $j$ for the $i$-th example. The prediction for the $i$-th point is the class with the highest score, i.e. $\text{argmax}_{j\in[c]}F^{*}_{ij}$.
@@ 3.1 Algorithm Families
We consider three parametric families, which we describe below.
@@
Here, $S$ is the symmetrically normalized adjacency matrix. This score matrix $F^{*}_{\alpha}$ corresponds to minimizing the following objective function $\mathcal{Q}(F)=\frac{1}{2}(\sum_{i,j=1}^{n}W_{ij}\|\frac{1}{\sqrt{d_{i}}}F_{i}-\frac{1}{\sqrt{d_{j}}}F_{j}\|^{2}+\frac{1-\alpha}{\alpha}\sum_{i=1}^{n}\|F_{i}-Y_{i}\|^{2}).$ The first term of $Q(F)$ measures the local consistency, i.e., the prediction between nearby points should be similar. The second term measures the global consistency, i.e., consistency with respect to its original label. Therefore, the parameter $\alpha\in(0,1)$ induces a trade-off between the local and the global consistency. We denote this family as $\mathcal{F}_{\alpha}$, and the corresponding 0-1 losses as $\mathcal{H}_{\alpha}$.
@@
Here we consider a new algorithm family which we name Normalized Adjacency Matrix Based Family. This class of algorithm is parameterized by $\delta\in[0,1]$. The scoring matrix $F^{*}_{\delta}$ is
$F^{*}_{\delta}=(I-c\cdot S)^{-1}Y,~{}~{}\text{where }S=D^{-\delta}WD^{\delta-1}.$
Here, $S$ is the (not necessarily symmetrically) normalized adjacency matrix and $c\in\mathbb{R}$ is a constant.
@@ 3.2 Pseudo-dimension Guarantees
The proofs of the above three theorems follow a similar approach. Here, we give an overview of the proof idea. The full proofs are in Appendix A.
@@ 4 GNN Families and Generalization Guarantees
In Section 4.2, we introduce a novel architecture, which we call Graph Convolution and Attention Networks (GCAN), that uses a hyperparameter $\eta\in[0,1]$ to interpolate two popular GNNs: the graph convolutional neural networks (GCN) and graph attention neural networks (GAT). GCAN selects the optimal model for specific datasets: $\eta=0$ corresponds to GCN, $\eta=1$ to GAT, and intermediate values explore hybrid architectures that may outperform both. We also establish a Rademacher complexity bound for the GCAN family.
@@ 5 Experiments
The results are shown in Figure 2. Note that GCN outperforms GAT on some datasets (e.g. CORA, Actor) and GAT performs better on others (e.g. CIFAR10, see also Dwivedi et al. (2023)). With GCAN, we can achieve the best performance on each dataset–in Figure 2, GCAN consistently achieves higher or comparable accuracy compared to both GAT and GCN across all datasets. Notably, GCAN demonstrates significant improvements in CIFAR10 and CORA, highlighting its effectiveness in these scenarios.
@@ Appendix C Proofs in Section 4
$=$ $\|\left(\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)+\sum_{j=1}^{n}\frac{w_{ij}T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right)$
$-\left(\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})+\sum_{j=1}^{n}\frac{w_{ij}T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right)\|$
$\leq$ $\|\left(\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})\right)\|$
$+\sum_{j=1}^{n}\left(\|w_{ij}\|\|\left(\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right)\|\right)$ (by triangle inequality)
@@
$\leq$ $\left\|\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta,\theta)\right\|$
$+\left\|\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})\right\|$ (by triangle inequality)
$\leq$ $\left\|\frac{\beta}{d_{i}+\beta}-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}\right\|\left\|T_{L-1,i}(\beta,\theta)\right\|+\left\|\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}\right\|\Delta_{L-1,i}$ (by Cauchy-Schwarz inequality)
@@
$\leq$ $\left\|\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|$
$+\left\|\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}-\frac{T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|$ (by triangle inequality)
$\leq$ $\left\|\frac{1}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{1}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|\|T_{L-1,j}(\beta,\theta)\|$
$+\left\|\frac{1}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|\Delta_{L-1,i}$ (Cauchy-Schwarz inequality)
@@
$=$ $\left\|\frac{(d_{i}+\beta)(d_{j}+\beta)-(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}[\sqrt{(d_{i}+\beta)(d_{j}+\beta)}+\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}]}\right\|$
$\leq$ $\left\|\frac{(d_{i}+d_{j}+\beta+\beta^{\prime})(\beta-\beta^{\prime})}{C_{dl}+\beta)(C_{dl}+\beta^{\prime})[(C_{dl}+\beta)+(C_{dl}+\beta^{\prime})]}\right\|$
$\leq$ $\frac{C_{dh}+1}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|$
@@
$+\sum_{i=1}^{n}\left(\|w_{ij}\|\left(\frac{(C_{dh}+1)\|T_{L-1,i}(\beta,\theta)\|}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|+\frac{1}{C_{dl}}\Delta_{L-1,i}\right)\right)$
@@
$+d_{i}\left(\frac{(C_{dh}+1)\|T_{L-1,i}(\beta,\theta)\|}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|+\frac{1}{C_{dl}}\Delta_{L-1,i}\right)$
@@
$\|T_{L-1,i}(\beta,\theta)\|$ $=\|\frac{\beta}{d_{i}+\beta}T_{L-2,i}(\beta,\theta)+\sum_{j=1}^{n}\frac{w_{ij}T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\|$
$\leq\|\frac{\beta}{d_{i}+\beta}T_{L-2,i}(\beta,\theta)\|+\sum_{j=1}^{n}\|\frac{w_{ij}T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\|$ (by triangle inequality)
$\leq\frac{\beta}{d_{i}+\beta}\|T_{L-2,i}(\beta,\theta)\|+\sum_{j=1}^{n}\|w_{ij}\|\|\frac{T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\|$ (by Cauchy-Schwarz)
$\leq\frac{\beta}{d_{i}+\beta}\|T_{L-2,i}(\beta,\theta)\|+C_{dh}\max_{j}\|\frac{T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\|$
@@
$\left\|\left(\eta\cdot e_{ij}^{L-1}Uh_{j}^{L-1}\right)-\left(\eta^{\prime}\cdot e_{ij}^{\prime(L-1)}U^{\prime}h_{j}^{\prime(L-1)}\right)\right\|$
$\leq$ $C_{U}\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\cdot|\eta-\eta^{\prime}|+C_{U}\bar{h}_{j}^{L-1}\cdot\left|e_{ij}^{L-1}-e_{ij}^{\prime(L-1)}\right|$
$+\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\left\|U-U^{\prime}\right\|+C_{U}\bar{e}_{ij}^{L-1}\left\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\right\|$
@@
$\|(1-\eta)\cdot\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-(1-\eta^{\prime})\cdot\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\|$
$\leq$ $\|\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\|+\|\eta\cdot\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-\eta^{\prime}\cdot\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\|$ (by triangle inequality)
$\leq$ $\frac{1}{C_{dl}}\|Uh_{i}^{L-1}-U^{\prime}h_{i}^{\prime(L-1)}\|+\frac{1}{C_{dl}}\|\eta\cdot Uh_{i}^{L-1}-\eta^{\prime}\cdot U^{\prime}h_{i}^{\prime(L-1)}\|$
$\leq$ $\frac{1}{C_{dl}}\left(C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+\bar{h}_{i}^{L-1}\|U-U^{\prime}\|\right)$
$+\frac{1}{C_{dl}}\left(C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+C_{U}\bar{h}_{i}^{L-1}|\eta-\eta^{\prime}|\right)$ (using Lemma C.5)
$=$ $\frac{1}{C_{dl}}\left(2C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+2\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+C_{U}\bar{h}_{i}^{L-1}|\eta-\eta^{\prime}|\right).$
@@
$\Delta_{i}^{L}\leq$ $\sum_{j\in\mathcal{N}_{i}}\Big{(}C_{U}\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\cdot\left|\eta-\eta^{\prime}\right|+C_{U}\bar{h}_{j}^{L-1}\cdot\left|e_{ij}^{L-1}-e_{ij}^{\prime(L-1)}\right|$
$+\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\left\|U-U^{\prime}\right\|+C_{U}\bar{e}_{ij}^{L-1}\left\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\right\|$
@@
$\leq$ $C_{U}\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)|\eta-\eta^{\prime}|+rC_{U}\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)+\left(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1}\right)\|U-U^{\prime}\|$
@@
$\leq C_{U}\sum_{j\in\mathcal{N}_{i}}\left|\eta\cdot e_{ij}^{\ell}+(1-\eta)\cdot\frac{1}{\sqrt{d_{i}d_{j}}}\right|\left\|h_{j}^{\ell}\right\|$
$\leq rC_{U}\max\left(1,\frac{1}{C_{dl}}\right)\left(\max_{j\in\mathcal{N}_{i}}\left\|h_{j}^{\ell-1}\right\|\right)$
@@
$k_{3}$ $=\left(1+\frac{2r}{C_{dl}}\right)r^{L-1}C_{U}^{L}C_{z}\max\left(1,\frac{1}{C_{dl}}\right)^{L-1}$

### v2

@@ 3 Label Propagation-based Families and Generalization Guarantees
In this section, we consider three parametric families of label propagation-based algorithms, the classical type of algorithms for semi-supervised learning. Label propagation algorithms output a soft-label score $F^{*}\in\mathbb{R}^{n\times c}$, where the $(i,j)$-th entry of $F^{*}$ represents the score of class $j$ for the $i$-th sample. The prediction for the $i$-th sample is the class with the highest score, i.e. $\text{argmax}_{j\in[c]}F^{*}_{ij}$.
@@ 3.1 Algorithm Families
We consider three parametric families described below.
@@
Here, $S$ is the symmetrically normalized adjacency matrix. This score matrix $F^{*}_{\alpha}$ corresponds to minimizing the following objective function $\mathcal{Q}(F)=\frac{1}{2}(\sum_{i,j=1}^{n}W_{ij}\|\frac{1}{\sqrt{d_{i}}}F_{i}-\frac{1}{\sqrt{d_{j}}}F_{j}\|^{2}+\frac{1-\alpha}{\alpha}\sum_{i=1}^{n}\|F_{i}-Y_{i}\|^{2}).$ The first term of $Q(F)$ measures the local consistency, i.e., the prediction between nearby points should be similar. The second term measures the global consistency, i.e., consistency to its original label. Therefore, the parameter $\alpha\in(0,1)$ induces a trade-off between the local and the global consistency. We denote this family as $\mathcal{F}_{\alpha}$, and the 0-1 losses as $\mathcal{H}_{\alpha}$.
@@
Here we consider an algorithm family (Avrachenkov et al., 2012). This class of algorithm is parameterized by $\delta\in[0,1]$. The scoring matrix $F^{*}_{\delta}$ is
$F^{*}_{\delta}=(I-c\cdot S)^{-1}Y,~{}~{}\text{where}~{}~{}S=D^{-\delta}WD^{\delta-1}.$
Here, $S$ is the (not symmetrically) normalized adjacency matrix and $c\in\mathbb{R}$ is a constant.
@@ 3.2 Pseudo-dimension Guarantees
Similar to Local and Global Consistency, we give a tight $\Theta(\log n)$ bound on the pseudo-dimension of the Smoothing-based family of Delalleau et al. (2005).
@@
Finally, and perhaps more surprisingly, we give the same $\Theta(\log n)$ bound on the pseudo-dimension of the normalized adjacency-matrix based family.
@@
The proofs of the above three theorems follow a similar template. Here, we give an overview of the proof idea. The full proof is in Appendix A.
@@ 4 GNN Families and Generalization Guarantees
In Section 4.2, we introduce a novel architecture, which we call GCAN, that uses a hyperparameter $\eta\in[0,1]$ to interpolate two popular GNNs: the graph convolutional neural networks (GCN) and graph attention neural networks (GAT). GCAN selects the optimal model for specific datasets: $\eta=0$ corresponds to GCN, $\eta=1$ to GAT, and intermediate values explore hybrid architectures that may outperform both. We also establish a Rademacher complexity bound for the GCAN family.
@@ 5 Experiments
We also compare our backpropagation-based approach with Bayesian Optimization (see e.g. Frazier (2018)). Using the same 20 training sub-graphs, we perform Bayesian Optimization to select the hyperparameter $\eta$, ensuring both methods use an equal number of forward passes. The selected $\eta$ is then evaluated on a separate set of 20 test sub-graphs from the same dataset.
The results on the test set are shown in Figure 2. Note that GCN outperforms GAT on some datasets (e.g. CORA, CoAuthorCS) and GCN performs better on others (e.g. CIFAR10, see also Dwivedi et al. (2023)). With GCAN, we can achieve the best performance on most datasets. Indeed, as seen in Figure 2, GCAN consistently achieves higher or comparable accuracy compared to both GAT and GCN across all datasets. Notably, GCAN demonstrates significant improvements in CIFAR10 and CoAuthorCS, highlighting its effectiveness in these scenarios. Also, comparing backpropagation with Bayesian Optimization, backpropagation achieves better performance on more datasets (e.g.  CIFAR10, CoAuthorCS, Actor), but Bayesian Optimization is more effective in certain datasets (e.g.  CORA, AmazonPhotos).
@@ 6 Conclusion
An interesting direction for further investigation involves improving computational efficiency. Sharma and Jones (2023) introduced techniques for approximating loss functions to reduce the cost of tuning graph kernel hyperparameters. It would be worthwhile to explore whether their methods can be adapted to our setting to alleviate computational burdens.
Another natural extension is the tuning of multiple hyperparameters. Although this increases analytical complexity, we anticipate that our techniques remain applicable. For GNN-based algorithms, Rademacher Complexity may still provide a suitable capacity measure. Our approach would aim to bound the variation in predicted values as hyperparameters change and then apply a covering argument. For label propagation methods, we would instead analyze how the scoring matrix evolves with respect to hyperparameter shifts. These directions offer a compelling foundation for extending our techniques to more complex tuning scenarios.
Acknowledgments
We thank Nina Balcan and anonymous reviewers for helpful comments. This work was supported in part by the National Science Foundation under grants ECCS-2216899 and ECCS-2216970.
@@ Appendix C Proofs in Section 4
$=\Bigg{\|}\left(\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)+\sum_{j=1}^{n}\frac{w_{ij}T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right)$
$\qquad-\left(\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})+\sum_{j=1}^{n}\frac{w_{ij}T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right)\Bigg{\|}$
$\leq\left\|\left(\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})\right)\right\|$
$\qquad+\sum_{j=1}^{n}\left(\|w_{ij}\|\left\|\left(\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right)\right\|\right)$ (by triangle inequality)
@@
$\qquad\leq\left\|\frac{\beta}{d_{i}+\beta}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta,\theta)\right\|$
$\qquad\qquad+\left\|\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta,\theta)-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}T_{L-1,i}(\beta^{\prime},\theta^{\prime})\right\|$ (by triangle inequality)
$\qquad\leq\left\|\frac{\beta}{d_{i}+\beta}-\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}\right\|\left\|T_{L-1,i}(\beta,\theta)\right\|+\left\|\frac{\beta^{\prime}}{d_{i}+\beta^{\prime}}\right\|\Delta_{L-1,i}$ (by Cauchy-Schwarz inequality)
@@
$\qquad\leq\left\|\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|$
$\qquad\qquad+\left\|\frac{T_{L-1,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}-\frac{T_{L-1,j}(\beta^{\prime},\theta^{\prime})}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|$ (by triangle inequality)
$\qquad\leq\left\|\frac{1}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}-\frac{1}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|\|T_{L-1,j}(\beta,\theta)\|$
$\qquad\qquad+\left\|\frac{1}{\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}}\right\|\Delta_{L-1,i}$ (Cauchy-Schwarz inequality)
@@
$\qquad=\left\|\frac{(d_{i}+\beta)(d_{j}+\beta)-(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}[\sqrt{(d_{i}+\beta)(d_{j}+\beta)}+\sqrt{(d_{i}+\beta^{\prime})(d_{j}+\beta^{\prime})}]}\right\|$
$\qquad\leq\left\|\frac{(d_{i}+d_{j}+\beta+\beta^{\prime})(\beta-\beta^{\prime})}{C_{dl}+\beta)(C_{dl}+\beta^{\prime})[(C_{dl}+\beta)+(C_{dl}+\beta^{\prime})]}\right\|$
$\qquad\leq\frac{C_{dh}+1}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|$
@@
$\qquad+\sum_{i=1}^{n}\left(\|w_{ij}\|\left(\frac{(C_{dh}+1)\|T_{L-1,i}(\beta,\theta)\|}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|+\frac{1}{C_{dl}}\Delta_{L-1,i}\right)\right)$
@@
$\qquad+d_{i}\left(\frac{(C_{dh}+1)\|T_{L-1,i}(\beta,\theta)\|}{C_{dl}^{3}}\|\beta-\beta^{\prime}\|+\frac{1}{C_{dl}}\Delta_{L-1,i}\right)$
@@
$\|T_{L-1,i}(\beta,\theta)\|$ $=\left\|\frac{\beta}{d_{i}+\beta}T_{L-2,i}(\beta,\theta)+\sum_{j=1}^{n}\frac{w_{ij}T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right\|$
$\leq\left\|\frac{\beta}{d_{i}+\beta}T_{L-2,i}(\beta,\theta)\right\|+\sum_{j=1}^{n}\left\|\frac{w_{ij}T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right\|$ (by triangle inequality)
$\leq\frac{\beta}{d_{i}+\beta}\|T_{L-2,i}(\beta,\theta)\|+\sum_{j=1}^{n}\|w_{ij}\|\left\|\frac{T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right\|$ (by Cauchy-Schwarz)
$\leq\frac{\beta}{d_{i}+\beta}\|T_{L-2,i}(\beta,\theta)\|+C_{dh}\max_{j}\left\|\frac{T_{L-2,j}(\beta,\theta)}{\sqrt{(d_{i}+\beta)(d_{j}+\beta)}}\right\|$
@@
$\left\|(\eta\cdot e_{ij}^{L-1}Uh_{j}^{L-1})-(\eta^{\prime}\cdot e_{ij}^{\prime(L-1)}U^{\prime}h_{j}^{\prime(L-1)})\right\|$
$\qquad\leq C_{U}\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\cdot|\eta-\eta^{\prime}|+C_{U}\bar{h}_{j}^{L-1}\cdot\left|e_{ij}^{L-1}-e_{ij}^{\prime(L-1)}\right|$
$\qquad\qquad+\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\left\|U-U^{\prime}\right\|+C_{U}\bar{e}_{ij}^{L-1}\left\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\right\|$
@@
$\left\|(1-\eta)\cdot\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-(1-\eta^{\prime})\cdot\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\right\|$
$\qquad\leq\left\|\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\right\|+\left\|\eta\cdot\frac{1}{\sqrt{d_{i}d_{j}}}Uh_{i}^{L-1}-\eta^{\prime}\cdot\frac{1}{\sqrt{d_{i}d_{j}}}U^{\prime}h_{i}^{\prime(L-1)}\right\|$ (by triangle inequality)
$\qquad\leq\frac{1}{C_{dl}}\|Uh_{i}^{L-1}-U^{\prime}h_{i}^{\prime(L-1)}\|+\frac{1}{C_{dl}}\|\eta\cdot Uh_{i}^{L-1}-\eta^{\prime}\cdot U^{\prime}h_{i}^{\prime(L-1)}\|$
$\qquad\leq\frac{1}{C_{dl}}\left(C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+\bar{h}_{i}^{L-1}\|U-U^{\prime}\|\right)$
$\qquad\qquad+\frac{1}{C_{dl}}\left(C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+C_{U}\bar{h}_{i}^{L-1}|\eta-\eta^{\prime}|\right)$ (using Lemma C.5)
$\qquad=\frac{1}{C_{dl}}\left(2C_{U}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|+2\bar{h}_{i}^{L-1}\|U-U^{\prime}\|+C_{U}\bar{h}_{i}^{L-1}|\eta-\eta^{\prime}|\right).$
@@
$\Delta_{i}^{L}\leq$ $\sum_{j\in\mathcal{N}_{i}}\Big{(}C_{U}\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\cdot|\eta-\eta^{\prime}|+C_{U}\bar{h}_{j}^{L-1}\cdot|e_{ij}^{L-1}-e_{ij}^{\prime(L-1)}|$
$+\bar{e}_{ij}^{L-1}\bar{h}_{j}^{L-1}\|U-U^{\prime}\|+C_{U}\bar{e}_{ij}^{L-1}\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\|$
@@
$\leq$ $C_{U}(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})|\eta-\eta^{\prime}|+rC_{U}(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})+(\max_{j\in\mathcal{N}_{i}}\bar{h}_{j}^{L-1})\|U-U^{\prime}\|$
$+C_{U}(\max_{j\in\mathcal{N}_{i}}\|h_{j}^{L-1}-h_{j}^{\prime(L-1)}\|)+\frac{2rC_{U}}{C_{dl}}\|h_{i}^{L-1}-h_{i}^{\prime(L-1)}\|$
@@
$\leq C_{U}\sum_{j\in\mathcal{N}_{i}}\left|\eta\cdot e_{ij}^{\ell}+(1-\eta)\cdot\frac{1}{\sqrt{d_{i}d_{j}}}\right|\|h_{j}^{\ell}\|$
$\leq rC_{U}\max\left(1,\frac{1}{C_{dl}}\right)\left(\max_{j\in\mathcal{N}_{i}}\|h_{j}^{\ell-1}\|\right)$
@@
$k_{3}$ $=\left(1+\frac{2r}{C_{dl}}\right)r^{L-1}C_{U}^{L}C_{z}\max(1,\frac{1}{C_{dl}})^{L-1}$